OpenStudy (korosh23):

Sometimes a function has no name, and we see something like: y = x^2 Explain what is the input, and what is the out put?

OpenStudy (anonymous):

x is the input and y is the output

OpenStudy (anonymous):

same thing as a regular function

OpenStudy (korosh23):

How do you know y is the out put?

OpenStudy (korosh23):

@dtan5457 do you know?

OpenStudy (korosh23):

@praneet_r

OpenStudy (dtan5457):

Well I'm not too sure what your looking for here. y=x^2 to get y, you must input a x value.. you can't really do it the other way around

OpenStudy (korosh23):

what is our out put ?

OpenStudy (dtan5457):

y?

OpenStudy (korosh23):

I mathisfun.com it said the same thing, but it also gave a picture that is confusing

OpenStudy (dtan5457):

Oh that means let's say f(x)=x^2 f(5)=x^2 you input is 5, and your output is (5)^2 but that's basically the same thing lol i wouldn't worry to much about that

OpenStudy (korosh23):

Ok thank you lol